Suppose the market for landscape workers in Kitchener, Ontario, can be described by the following labour demand and supply equations: LD = 550 - 5 w and LS = 100 + 30 w.

Calculate the equilibrium wage and employment in this market.

equilibrium wage  

equilibrium employment  

Suppose the local government passes a bill requiring all employers in this industry to provide bottled water to each worker to protect them from hot weather. This has been estimated to cost employers $ 1.5 per worker per hour. This way, workers do not need to purchase water by themselves, so that the bill effectively increases the wage by $ 1.5 per hour.

Calculate the new equilibrium wage and employment.

equilibrium wage  

equilibrium employment  
Answer #1

i)

LD=550-5w

LS=100+30w

Set LD=LS for equilibrium

550-5w=100+30w

35w=450

w=450/35 =$12.8571

LD=550-5w=550-5*12.8571=485.71

LS=100+30*12.8571=485.71

Equilibrium wage=$12.86

Equilibrium employment=485.71

ii)

In this case supply curve remains unchanged while demand curve changes as

LD'=550-5*(w+1.50)=550-5w-7.50=542.50-5w

LS=100+30w

Set LD'=LS for equilibrium

542.50-5w=100+30w

35w=442.50

w=12.64286

LD'=542.50-5*12.64286=479.2857

LS=100+30*12.6429=479.2858

So, New equilibrium wage=$12.64 per hour

New equilibrium employment=479.29